The speed of bus is 25 km /hr to 70 km/hr in 5 sec find acceleration

Grade:9

1 Answers

poonam
6 years ago
Given in question initial velocity (u ) =25km/hr =6.94m/sec Final velocity (v)= 70km/hr = 19.44m/sec Time(t) = 5sec We know that acceleration (a)= v-u/t =19.44-6.94/5 = 2.5 therfore acceleration = 2.5m/sec^2